OpenStudy (anonymous):

Find the Vertex ( X and Y coordinate), the line of symmetry, and determine if the value is minimum or maximum f(x)=-2x^2+2x+6

OpenStudy (saifoo.khan):

Let's use a shortcut, -b/2a to find the vertex. -(2)/2(-2) -2/-4 1/2 x cordinate = 1/2 -2x^2+2x+6 -2(1/2)^2 +2(1/2) + 6 13/2

OpenStudy (saifoo.khan):

Therefore, x-cordinate = 1/2 y-cordinate = 13/2

OpenStudy (saifoo.khan):

Line of symm = x =1/2 Since the value of a is negative, so it will be a happy face, a minimum point.

OpenStudy (saifoo.khan):